If

Consumption function C= 100+ 0.5Yd

TaxT= 100

InvestmentI= 200

ExportX= 120

Government Expenditure G= 150

ImportM= 0.03Y

Find the following:

  1. The equilibrium level of income.
  2. The next exports.
1
Expert's answer
2022-03-07T11:02:41-0500
  1. The equilibrium level of income.

"Y=C+I+G+X-M"


"C=100+0.5Yd"


"=100+0.5(Y-T)"


"=100+0.5(Y-0.03Y)"


"Y=100+0.5(Y-0.03Y)+200+150+120-0.03Y"


"Y=100+0.5Y\u22120.015Y)+200+150+120\u22120.03Y"


"Y=570+0.455Y"


"Y-0.455Y=570"


"0.545Y=570"


"Y=1,045.87"


2). The next exports.


"The \space value \space of \space exports (X) -value \space of \space imports (M)"


"X-M"


"=120-(0.03\times 1,045.87)"


"=120-33.38"


"= 88.62"